# Neural-Augmented Kelvinlet for Real-Time Soft Tissue Deformation Modeling

### v2

@@ Abstract
Accurate and efficient modeling of soft-tissue interactions is fundamental for advancing surgical simulation, surgical robotics, and model-based surgical automation. To achieve real-time latency, classical Finite Element Method (FEM) solvers are often replaced with neural approximations; however, naively training such models in a fully data-driven manner without incorporating physical priors frequently leads to poor generalization and physically implausible predictions. We present a novel physics-informed neural simulation framework that enables real-time prediction of soft-tissue deformations under complex single- and multi-grasper interactions. Our approach integrates Kelvinlet-based analytical priors with large-scale FEM data, capturing both linear and nonlinear tissue responses. This hybrid design improves predictive accuracy and physical plausibility across diverse neural architectures while maintaining the low-latency performance required for interactive applications. We validate our method on challenging surgical manipulation tasks involving standard laparoscopic grasping tools, demonstrating substantial improvements in deformation fidelity and temporal stability over existing baselines. These results establish Kelvinlet-augmented learning as a principled and computationally efficient paradigm for real-time, physics-aware soft-tissue simulation in surgical AI. Our code and data is available at: https://github.com/Anon92373/Neural-Kelvinlet.
@@
Soft tissue deformation modeling is essential in computational biomechanics for surgical simulation, image registration, and robotic-assisted interventions. Accurate tissue simulation enables surgical planning, image-guided predictions of soft-tissue changes, and precise haptic feedback for robotics. Traditional approaches rely on physics-based methods grounded in continuum mechanics, such as the Finite Element Method (FEM), Boundary Element Method (BEM), and Tensor Mass Model (TMM). While highly accurate and physically realistic, these methods are often computationally prohibitive for real-time applications [2, 34]. To reduce computational cost, approximate methods such as mass-spring systems [23], mesh-free methods [3], and particle-based methods [21] are frequently used. However, these approximations often sacrifice realism, creating a significant simulation-to-reality gap that limits their practicality in high-fidelity applications.
@@
Recent advances in deep learning and neural operators have enabled fully data-driven methods that approximate soft tissue deformations with a single forward pass of a learned model. Convolutional neural networks (CNNs), graph neural networks (GNNs), and Transformers [26] have demonstrated the ability to capture complex deformation patterns from simulated or real-world data, enabling real-time tissue simulation. Despite these advances, fully data-driven approaches face two fundamental limitations: (a) limited generalization to unseen force distributions and boundary conditions, which results in unreliable extrapolation, and (b) a tendency to produce physically inconsistent deformations, particularly when trained on datasets of limited size or diversity. These shortcomings highlight the need for hybrid approaches that combine the expressiveness of neural models with the robustness of physics-based priors.
@@
To address these challenges, considerable efforts have focused on integrating physics priors into data-driven machine learning models [17]. Broadly, these methods make learning algorithms physics-informed by introducing: (a) observational biases, such as data augmentations that embed physical constraints; (b) inductive biases in the model architecture, for example, using group-invariant or equivariant neural networks [5]; and (c) learning biases through the design of loss functions that explicitly enforce physical laws. Despite this progress, achieving the necessary low latency for real-time simulation while maintaining a high-performing physics-informed model with appropriately chosen biases remains a challenging and largely open problem.
@@
In this paper, we model soft tissue using a volumetric mesh and address the problem of predicting the displacement of the entire mesh given the observed displacements at a limited set of interaction points on its surface. This is analogous to determining how the entire soft tissue will deform when graspers engage multiple surface locations and are moved to specified target positions. While several prior studies have explored similar setup and demonstrated that neural networks can approximate the solutions produced by high-fidelity physics-based solvers such as FEM [27], these approaches often lack strong physical priors and can suffer from reduced generalization and physical inconsistency. To overcome these limitations, we propose a physics-informed neural framework that integrates analytical elastic priors to enhance both accuracy and physical plausibility.
@@
We introduce a physics-informed neural framework that leverages Kelvinlets [8], i.e., the closed-form Green’s function solutions to the Navier–Cauchy equations, as an analytically tractable and computationally efficient prior. Kelvinlets provide a fast, low-order approximation of elastic deformation that captures the fundamental physics of linear elasticity. We build on this prior through two complementary hybrid strategies: (a) residual learning: using Kelvinlets to generate a baseline displacement field and training a neural network to model the residual error; and (b) physics regularization: incorporating Kelvinlets as a soft constraint within the learning objective of a neural network that predicts the full deformation field. By explicitly coupling the expressiveness of neural networks with the efficiency and physical grounding of Kelvinlets, our approach achieves improved accuracy, generalization, and physical plausibility for soft tissue deformation modeling.
@@ 1 Related Work
Classical approaches based on the Finite Element Method (FEM) provide high-fidelity simulations of soft tissue deformation but are often computationally prohibitive for real-time surgical applications. To address this limitation, model reduction techniques such as Proper Orthogonal Decomposition (POD) have been proposed to accelerate FEM without sacrificing physical realism. For example, Lau et al. [20] employ a reduced-order linear FEM with modal warping to achieve sub-15 ms inference for liver deformation, enabling intraoperative use. Complementarily, Hu and Desai [14] characterize nonlinear liver material behavior through indentation experiments and validate Local Effective Moduli (LEM) estimates using hyperelastic FEM models, laying foundational work for realistic soft-tissue mechanics under large strains.
@@
More recently, physics-informed neural networks (PINNs) have emerged as a data-efficient alternative that integrates governing equations such as elasticity or dynamics directly into the learning objective. Hu et al. [15] introduce a PINN that encodes static linear elasticity to predict 3D stress and strain fields of liver tissue under surgical tool interaction, achieving millisecond-scale inference while retaining physical plausibility. Extending this paradigm, Nguyen-Le et al. [24] propose a PINODE framework that augments neural dynamics with mass-spring priors, enabling extrapolation from sparse motion data. These methods improve interpretability and generalization but often require complex training and remain constrained by solver latency or limited scalability.
@@
In contrast, purely data-driven surrogate models have been proposed to directly regress deformation fields from control variables. Ke et al. [6] train a feedforward network to predict liver surface displacements conditioned on retractor pose, achieving sub-millisecond inference. While fast and simple, such models lack physical constraints, which can reduce reliability and robustness in unseen conditions.
@@
On the learning-based front, mesh-based models such as PhysGNN leverage FEM supervision to train graph neural networks for real-time deformation prediction in neurosurgery [33]. PEGNN [32] extends this with contact-aware hybrid graph reasoning, enabling accurate predictions across surgical and manipulation scenarios. Meanwhile, SeeSaw [9] employs stereo video and self-supervised learning to estimate sparse 3D deformations using a GNN encoder. Zhao et al. [12] introduce a diffeomorphic mapping framework that enforces cycle consistency, capturing invertible tissue motion. These approaches show strong empirical performance but may suffer from limited interpretability or reliance on dense supervision.
@@
Neural architectures such as U-Net [22], Set Attention Block (SAB) [19], and GraphGPS [29] offer modular and scalable backbones for modeling deformation fields across grid-based, set-based, and graph-structured data. While effective in capturing spatial patterns and long-range dependencies, these models are generally agnostic to the underlying physics of soft tissue behavior and typically lack inductive biases grounded in biomechanics, such as force equilibrium, boundary conditions, or constitutive material priors.
@@
2 Large-Scale FEM Data
@@
Neural network–based solutions require training on large-scale finite element (FE) simulation datasets. In this work, we focus on laparoscopic cholecystectomy and construct an FE model from an abdominal CT image and its segmentation masks, converting them into a 3D linear tetrahedral mesh with 10,400 nodes and 44,045 tetrahedra using gmsh [11]. As part of our contributions, we release this large-scale dataset; the details of its creation are provided below.
The bare area of the liver, where it contacts the inferior vena cava and diaphragm, and the distal ends of the ducts and veins are fixed (non-deformable). Tool interaction regions are selected to simulate the initial surgical step of lifting the liver lobes to expose the gallbladder and cystic ducts. These regions are identified by clustering nodes along the liver ridge, based on surgical video analysis and input from our surgical collaborator. FE simulations are performed using the open-source getFEM package [30], applying tool-imposed displacements at those predefined clustered regions. This pipeline is illustrated in Figure 1.
@@
For our nonlinear simulations, we use a two-parameter Mooney-Rivlin model [31], where the strain energy density can be written as functions of the Cauchy-Green deformation tensor invariants ($\boldsymbol{I_{i}(C)}$, with $i=1,2,3$) [13]:
@@
where $C_{10}=1.62$ kPa and $C_{01}=1.97$ kPa are constitutive constants derived from ex-vivo porcine liver experiments [4].
@@
In these linear simulations, we assign a Poisson’s ratio of 0.45 for all organs [7]. The shear modulus values are also set as follows: 0.03 kPa for the fascia [1], 0.34 kPa for the gallbladder [25], 0.69 kPa for the gallbladder wall and liver [18], and 1.07 kPa for the remaining organs [10].
2.0.1 Training Distribution, $p$.
@@
2.0.2 Distribution for Multiple Graspers.
@@
3 Neural-Augmented Kelvinlet
@@
3.1 Navier-Cauchy Eqs and Green’s Function Solution
@@
where $\Gamma_{i}(\mathbf{x})\in\mathbb{R}^{3\times 3}$. Note that this formulation enables the total displacement function to incorporate cross communication between interaction points.
@@
3.2 Method 1: Kelvinlet-Based Residual Learning
@@
3.3 Method 2: Kelvinlets as Regularization
@@
Figure 2 illustrates the proposed residual- and regularization-based learning methods, along with the displacement norm visualizations for all methods on a representative sample.
4 Experiments
4.0.1 Data.
@@ 4 Experiments
4.0.2 Neural Architectures.
We first evaluate several deformation-specific baselines designed for soft tissue modeling. PhysGNN [33] is a mesh-based GNN trained on FEM simulations; we enhance it with a global attention layer after each graph convolution to mitigate the well-known oversmoothing associated with GNNs [35]. PEGNN [32] models contact dynamics via cross-graph attention; we retain only its soft-tissue encoder to isolate local geometric features. For SeeSaw [9], we reimplement the method and adapt it to a point-cloud-based simulation setup. From the diffeomorphic mapping approach [12], we extract only the deformation network to enable direct supervision. We also include a Transformer-based KAN [16], limited to single-step prediction for fair comparison. To assess architecture-agnostic effects of Kelvinlet priors, we additionally evaluate U-Net [22], SAB [19], GraphGPS [29], and PointNet [28]. U-Net is converted into a DeepSet by replacing convolutions with shared MLPs. SAB uses two self-attention layers and an MLP head. GraphGPS and PointNet are simplified by reducing their depth to prevent overfitting on high-resolution meshes.
@@
4.1 Results and Analysis
Table 1: Deformation Capture Mean (DCM) (%) on individual graspers in linear simulations for modified baselines and our method, with and without Kelvinlet priors.
Ours
Model Base +Res. +Reg.
PhysGNN[33] 90.12 (0.13) 91.77 (0.11) 92.65 (0.10)
SeeSaw[9] 89.24 (0.15) 91.02 (0.10) 91.83 (0.08)
Physics-Encoded GNN[32] 88.55 (0.14) 90.44 (0.09) 91.58 (0.12)
Transformer-KAN[16] 90.89 (0.12) 92.01 (0.08) 93.22 (0.07)
DiffoMorphic Deformation[12] 91.14 (0.10) 92.88 (0.11) 94.33 (0.09)
PointNet[28] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09)
U-Net[22] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11)
GraphGPS[29] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23)
SAB[19] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04)
Kelvinlet 65.24 (0) - -
Average Gain +1.23 +2.39
Minimum/Maximum Gain +0.43/1.89 +1.01/3.19
Table 2: Inference time (ms) on an NVIDIA RTX A6000 GPU for a single grasper with individual models in linear simulations, with and without Kelvinlet priors.
Inference time (ms) Ours
Model Base +Res. +Reg.
PhysGNN[33] 1.71 (0.11) 1.84 (0.13) 1.71 (0.11)
SeeSaw[9] 1.56 (0.09) 1.69 (0.08) 1.56 (0.09)
Physics-Encoded GNN[32] 1.65 (0.08) 1.79 (0.08) 1.65 (0.08)
Transformer-KAN[16] 2.22 (0.17) 2.32 (0.16) 2.22 (0.17)
DiffoMorphic Deformation[12] 2.16 (0.12) 2.35 (0.11) 2.16 (0.12)
PointNet[28] 1.16 (0.10) 1.23 (0.13) 1.16 (0.10)
U-Net[22] 1.55 (0.08) 1.64 (0.07) 1.55 (0.08)
GraphGPS[29] 2.60 (0.15) 2.78 (0.15) 2.60 (0.15)
SAB[19] 2.30 (0.02) 2.44 (0.21) 2.30 (0.20)
@@ 4.1 Results and Analysis
Table 3: Performance of various models on linear simulations with individual graspers, combined graspers, and multitask learning. DCM (%) is reported for single-task and multi-task settings with standard deviations.
Model Indiv. Graspers (DCM) Comb. Graspers (DCM) Multitask (DCM)
@@
U -Net[22] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11) 88.23 (0.08) 89.67 (0.09) 90.01 (0.08) 91.20 (0.09) 92.33 (0.10) 92.88 (0.08)
GraphGPS[29] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23) 87.33 (0.06) 87.92 (0.02) 88.01 (0.04) 88.35 (0.13) 88.91 (0.10) 89.45 (0.09)
PointNet[28] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09) 89.79 (0.07) 90.11 (0.07) 91.22 (0.07) 94.00 (0.08) 94.12 (0.07) 95.01 (0.06)
SAB[19] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04) 88.05 (0.09) 89.91 (0.11) 91.02 (0.04) 92.54 (0.06) 93.66 (0.07) 94.42 (0.06)
@@
Table 4: Performance of various models on nonlinear simulations with individual graspers, combined graspers, and multitask learning. DCM (%) is reported for single-task and multi-task settings with standard deviations.
Model Indiv. Graspers (DCM) Comb. Graspers (DCM) Multitask (DCM)
@@
U -Net[22] 80.56 (0.10) 84.22 (0.09) 84.88 (0.11) 76.91 (0.05) 79.83 (0.08) 82.10 (0.21) 78.10 (0.12) 80.45 (0.09) 81.33 (0.14)
GraphGPS[29] 77.52 (0.17) 80.63 (0.13) 82.11 (0.09) 73.72 (0.15) 77.92 (0.10) 80.66 (0.16) 76.24 (0.18) 79.31 (0.15) 80.42 (0.10)
PointNet[28] 81.25 (0.22) 84.65 (0.19) 85.05 (0.15) 79.56 (0.18) 82.66 (0.16) 83.33 (0.11) 83.25 (0.17) 84.75 (0.13) 85.21 (0.09)
SAB[19] 80.75 (0.19) 84.44 (0.07) 84.75 (0.10) 78.88 (0.11) 81.50 (0.15) 82.83 (0.16) 81.22 (0.10) 83.01 (0.08) 84.00 (0.07)

### v3

@@ Abstract
Accurate and efficient modeling of soft-tissue interactions is fundamental for advancing surgical simulation, surgical robotics, and model-based surgical automation. To achieve real-time latency, classical Finite Element Method (FEM) solvers are often replaced with neural approximations; however, naively training such models in a fully data-driven manner without incorporating physical priors frequently leads to poor generalization and physically implausible predictions. We present a novel physics-informed neural simulation framework that enables real-time prediction of soft-tissue deformations under complex single- and multi-grasper interactions. Our approach integrates Kelvinlet-based analytical priors with large-scale FEM data, capturing both linear and nonlinear tissue responses. This hybrid design improves predictive accuracy and physical plausibility across diverse neural architectures while maintaining the low-latency performance required for interactive applications. We validate our method on challenging surgical manipulation tasks involving standard laparoscopic grasping tools, demonstrating substantial improvements in deformation fidelity and temporal stability over existing baselines. These results establish Kelvinlet-augmented learning as a principled and computationally efficient paradigm for real-time, physics-aware soft-tissue simulation in surgical AI.
Code and datahttps://github.com/mint-vu/neural-augmented-kelvinlet-leap
Introduction
@@
Soft tissue deformation modeling is essential in computational biomechanics for surgical simulation, image registration, and robotic-assisted interventions. Accurate tissue simulation enables surgical planning, image-guided predictions of soft-tissue changes, and precise haptic feedback for robotics. Traditional approaches rely on physics-based methods grounded in continuum mechanics, such as the Finite Element Method (FEM), Boundary Element Method (BEM), and Tensor Mass Model (TMM). While highly accurate and physically realistic, these methods are often computationally prohibitive for real-time applications (Allard, Courtecuisse, and Faure 2012; Zhang, Zhong, and Gu 2019). To reduce computational cost, approximate methods such as mass-spring systems (Nedel and Thalmann 1998), mesh-free methods (Belytschko et al. 1996), and particle-based methods (Liu et al. 2015) are frequently used. However, these approximations often sacrifice realism, creating a significant simulation-to-reality gap that limits their practicality in high-fidelity applications.
@@
Recent advances in deep learning and neural operators have enabled fully data-driven methods that approximate soft tissue deformations with a single forward pass of a learned model. Convolutional neural networks (CNNs), graph neural networks (GNNs), and Transformers (Wu, Kazanzides, and Unberath 2020; Wu et al. 2021; Pfaff et al. 2021) have demonstrated the ability to capture complex deformation patterns from simulated or real-world data, enabling real-time tissue simulation. Despite these advances, fully data-driven approaches face two fundamental limitations: (a) limited generalization to unseen force distributions and boundary conditions, which results in unreliable extrapolation, and (b) a tendency to produce physically inconsistent deformations, particularly when trained on datasets of limited size or diversity. These shortcomings highlight the need for hybrid approaches that combine the expressiveness of neural models with the robustness of physics-based priors.
@@
To address these challenges, considerable efforts have focused on integrating physics priors into data-driven machine learning models (Karniadakis et al. 2021). Broadly, these methods make learning algorithms physics-informed by introducing: (a) observational biases, such as data augmentations that embed physical constraints; (b) inductive biases in the model architecture, for example, using group-invariant or equivariant neural networks (Bronstein et al. 2017); and (c) learning biases through the design of loss functions that explicitly enforce physical laws. Despite this progress, achieving the necessary low latency for real-time simulation while maintaining a high-performing physics-informed model with appropriately chosen biases remains a challenging and largely open problem.
@@
In this paper, we model soft tissue using a volumetric mesh and address the problem of predicting the displacement of the entire mesh given the observed displacements at a limited set of interaction points on its surface. This is analogous to determining how the entire soft tissue will deform when graspers engage multiple surface locations and are moved to specified target positions. While several prior studies have explored similar setup and demonstrated that neural networks can approximate the solutions produced by high-fidelity physics-based solvers such as FEM (Pfeiffer et al. 2019), these approaches often lack strong physical priors and can suffer from reduced generalization and physical inconsistency. To overcome these limitations, we propose a physics-informed neural framework that integrates analytical elastic priors to enhance both accuracy and physical plausibility.
@@
We introduce a physics-informed neural framework that leverages Kelvinlets (De Goes and James 2017), i.e., the closed-form Green’s function solutions to the Navier–Cauchy equations, as an analytically tractable and computationally efficient prior. Kelvinlets provide a fast, low-order approximation of elastic deformation that captures the fundamental physics of linear elasticity. We build on this prior through two complementary hybrid strategies: (a) residual learning: using Kelvinlets to generate a baseline displacement field and training a neural network to model the residual error; and (b) physics regularization: incorporating Kelvinlets as a soft constraint within the learning objective of a neural network that predicts the full deformation field. By explicitly coupling the expressiveness of neural networks with the efficiency and physical grounding of Kelvinlets, our approach achieves improved accuracy, generalization, and physical plausibility for soft tissue deformation modeling.
@@
Classical approaches based on the Finite Element Method (FEM) provide high-fidelity simulations of soft tissue deformation but are often computationally prohibitive for real-time surgical applications. To address this limitation, model reduction techniques such as Proper Orthogonal Decomposition (POD) have been proposed to accelerate FEM without sacrificing physical realism. For example, Lau et al. (Liu et al. 2025) employ a reduced-order linear FEM with modal warping to achieve sub-15 ms inference for liver deformation, enabling intraoperative use. Complementarily, Hu and Desai (Hu and Desai 2004) characterize nonlinear liver material behavior through indentation experiments and validate Local Effective Moduli (LEM) estimates using hyperelastic FEM models, laying foundational work for realistic soft-tissue mechanics under large strains.
@@
More recently, physics-informed neural networks (PINNs) have emerged as a data-efficient alternative that integrates governing equations such as elasticity or dynamics directly into the learning objective. Hu et al. (Hu et al. 2025) introduce a PINN that encodes static linear elasticity to predict 3D stress and strain fields of liver tissue under surgical tool interaction, achieving millisecond-scale inference while retaining physical plausibility. Extending this paradigm, Nguyen-Le et al. (Nguyen-Le, Ballit, and Dao 2025) propose a PINODE framework that augments neural dynamics with mass-spring priors, enabling extrapolation from sparse motion data. These methods improve interpretability and generalization but often require complex training and remain constrained by solver latency or limited scalability.
@@
In contrast, purely data-driven surrogate models have been proposed to directly regress deformation fields from control variables. Ke et al. (Camara et al. 2016) train a feedforward network to predict liver surface displacements conditioned on retractor pose, achieving sub-millisecond inference. While fast and simple, such models lack physical constraints, which can reduce reliability and robustness in unseen conditions.
@@
On the learning-based front, mesh-based models such as PhysGNN leverage FEM supervision to train graph neural networks for real-time deformation prediction in neurosurgery (Salehi and Giannacopoulos 2022). PEGNN (Saleh et al. 2024) extends this with contact-aware hybrid graph reasoning, enabling accurate predictions across surgical and manipulation scenarios. Meanwhile, SeeSaw (Docea et al. 2024) employs stereo video and self-supervised learning to estimate sparse 3D deformations using a GNN encoder. Zhao et al. (Gong et al. 2024) introduce a diffeomorphic mapping framework that enforces cycle consistency, capturing invertible tissue motion. These approaches show strong empirical performance but may suffer from limited interpretability or reliance on dense supervision.
@@
Neural architectures such as U-Net (Mendizabal, Márquez-Neila, and Cotin 2020), Set Attention Block (SAB) (Lee et al. 2019), and GraphGPS (Rampášek et al. 2022) offer modular and scalable backbones for modeling deformation fields across grid-based, set-based, and graph-structured data. While effective in capturing spatial patterns and long-range dependencies, these models are generally agnostic to the underlying physics of soft tissue behavior and typically lack inductive biases grounded in biomechanics, such as force equilibrium, boundary conditions, or constitutive material priors.
@@
Large-Scale FEM Data
@@
Neural network–based solutions require training on large-scale finite element (FE) simulation datasets. In this work, we focus on laparoscopic cholecystectomy and construct an FE model from an abdominal CT image and its segmentation masks, converting them into a 3D linear tetrahedral mesh with 10,400 nodes and 44,045 tetrahedra using gmsh (Geuzaine and Remacle 2009). As part of our contributions, we release this large-scale dataset; the details of its creation are provided below.
The bare area of the liver, where it contacts the inferior vena cava and diaphragm, and the distal ends of the ducts and veins are fixed (non-deformable). Tool interaction regions are selected to simulate the initial surgical step of lifting the liver lobes to expose the gallbladder and cystic ducts. These regions are identified by clustering nodes along the liver ridge, based on surgical video analysis and input from our surgical collaborator. FE simulations are performed using the open-source getFEM package (Renard and Poulios 2020), applying tool-imposed displacements at those predefined clustered regions. This pipeline is illustrated in Figure 1.
@@
For our nonlinear simulations, we use a two-parameter Mooney-Rivlin model (Rivlin 1948), where the strain energy density can be written as functions of the Cauchy-Green deformation tensor invariants ($\boldsymbol{I_{i}(C)}$, with $i=1,2,3$) (Hackett 2018):
@@
where $C_{10}=1.62$ kPa and $C_{01}=1.97$ kPa are constitutive constants derived from ex-vivo porcine liver experiments (Benson, Pereira, and Miga 2025).
@@
In these linear simulations, we assign a Poisson’s ratio of 0.45 for all organs (Chen et al. 1996). The shear modulus values are also set as follows: 0.03 kPa for the fascia (Akhmanova et al. 2015), 0.34 kPa for the gallbladder (Nisansala et al. 2016), 0.69 kPa for the gallbladder wall and liver (Kim et al. 2013), and 1.07 kPa for the remaining organs (George et al. 2018).
Training Distribution, $p$.
@@
Distribution for Multiple Graspers.
@@
Neural-Augmented Kelvinlet
@@
Navier-Cauchy Eqs and Green’s Function Solution
@@
where $\Gamma_{i}(\mathbf{x})\in\mathbb{R}^{3\times 3}$. and $r_{i,\varepsilon}(\mathbf{x}):=\sqrt{\|\mathbf{x}-\mathbf{x}_{i}\|_{2}^{2}+\varepsilon^{2}}.$ Note that this formulation enables the total displacement function to incorporate cross communication between interaction points.
@@
Method 1: Kelvinlet-Based Residual Learning
@@
Method 2: Kelvinlets as Regularization
@@
Figure  2 illustrates the proposed residual- and regularization-based learning methods, along with the displacement norm visualizations for all methods on a representative sample.
Experiments
Data.
@@
Neural Architectures.
We evaluate two groups of models: deformation-specific FEM surrogates and generic set/graph architectures. For the former, PhysGNN (Salehi and Giannacopoulos 2022) is a mesh-based GNN that we augment with a global attention layer to counter oversmoothing (Zhao and Akoglu 2020); PEGNN (Saleh et al. 2024) is used only through its soft-tissue encoder; SeeSaw (Docea et al. 2024) is reimplemented for our point-cloud setting; from the diffeomorphic framework of (Gong et al. 2024) we keep only the deformation network; and a Transformer-based KAN (Jiang et al. 2025) is restricted to single-step prediction. To study architecture-agnostic effects of Kelvinlet priors, we also use U-Net (Mendizabal, Márquez-Neila, and Cotin 2020), SAB (Lee et al. 2019), GraphGPS (Rampášek et al. 2022), and PointNet (Qi et al. 2016): U-Net is converted into a DeepSet-style model via shared pointwise MLPs, SAB uses two self-attention blocks with an MLP head, and GraphGPS and PointNet are shallow variants with reduced depth to avoid overfitting on high-resolution meshes.
@@
Results and Analysis
Model Base Ours
+Res. +Reg.
PhysGNN(Salehi and Giannacopoulos 2022) 1.71 (0.11) 1.84 (0.13) 1.71 (0.11)
SeeSaw(Docea et al. 2024) 1.56 (0.09) 1.69 (0.08) 1.56 (0.09)
Physics-Encoded GNN(Saleh et al. 2024) 1.65 (0.08) 1.79 (0.08) 1.65 (0.08)
Transformer-KAN(Jiang et al. 2025) 2.22 (0.17) 2.32 (0.16) 2.22 (0.17)
Diffeomorphic Deformation(Gong et al. 2024) 2.16 (0.12) 2.35 (0.11) 2.16 (0.12)
PointNet(Qi et al. 2016) 1.16 (0.10) 1.23 (0.13) 1.16 (0.10)
U-Net(Mendizabal, Márquez-Neila, and Cotin 2020) 1.55 (0.08) 1.64 (0.07) 1.55 (0.08)
GraphGPS(Rampášek et al. 2022) 2.60 (0.15) 2.78 (0.15) 2.60 (0.15)
SAB(Lee et al. 2019) 2.30 (0.02) 2.44 (0.21) 2.30 (0.20)
@@
Table 1: Inference time (ms) on an NVIDIA RTX A6000 GPU for a single grasper in linear simulations, with and without Kelvinlet priors. Values are mean (std).
Model Indiv. (DCM) Comb. (DCM) Multi (DCM)
@@
PhysGNN(Salehi and Giannacopoulos 2022) 90.12(0.13) 91.77(0.11) 92.65(0.10) 86.38(0.13) 88.10(0.11) 88.66(0.10) 89.55(0.13) 90.96(0.11) 91.54(0.10)
SeeSaw(Docea et al. 2024) 89.24(0.15) 91.02(0.10) 91.83(0.08) 85.50(0.15) 87.35(0.10) 87.84(0.08) 88.67(0.15) 90.21(0.10) 90.72(0.08)
Physics-Encoded GNN(Saleh et al. 2024) 88.55(0.14) 90.44(0.09) 91.58(0.12) 84.81(0.14) 86.77(0.09) 87.59(0.12) 87.98(0.14) 89.63(0.09) 90.47(0.12)
Transformer-KAN(Jiang et al. 2025) 90.89(0.12) 92.01(0.08) 93.22(0.07) 87.15(0.12) 88.34(0.08) 89.23(0.07) 90.32(0.12) 91.20(0.08) 92.11(0.07)
Diffeomorphic Deformation(Gong et al. 2024) 91.14(0.10) 92.88(0.11) 94.33(0.09) 87.40(0.10) 89.21(0.11) 90.34(0.09) 90.57(0.10) 92.07(0.11) 93.22(0.09)
PointNet(Qi et al. 2016) 94.23(0.11) 94.66(0.09) 96.54(0.09) 89.79(0.07) 90.11(0.07) 91.22(0.07) 94.00(0.08) 94.12(0.07) 95.01(0.06)
U-Net(Mendizabal, Márquez-Neila, and Cotin 2020) 92.11(0.12) 93.56(0.10) 94.04(0.11) 88.23(0.08) 89.67(0.09) 90.01(0.08) 91.20(0.09) 92.33(0.10) 92.88(0.08)
GraphGPS(Rampášek et al. 2022) 88.60(0.21) 89.45(0.15) 89.61(0.23) 87.33(0.06) 87.92(0.02) 88.01(0.04) 88.35(0.13) 88.91(0.10) 89.45(0.09)
SAB(Lee et al. 2019) 93.42(0.11) 94.60(0.04) 96.03(0.04) 88.05(0.09) 89.91(0.11) 91.02(0.04) 92.54(0.06) 93.66(0.07) 94.42(0.06)
@@
Table 2: Performance of various models on linear simulations with individual graspers, combined graspers, and multitask learning. DCM (%) is reported with standard deviations. “Base” denotes the original model, “+Res.” augments it with a Kelvinlet-based residual, and “+Reg.” applies a Kelvinlet-inspired regularizer.
Model Indiv. (DCM) Comb. (DCM) Multi (DCM)
@@
PhysGNN(Salehi and Giannacopoulos 2022) 78.05(0.13) 82.19(0.11) 82.79(0.10) 75.30(0.13) 79.18(0.11) 80.83(0.10) 77.73(0.13) 80.58(0.11) 81.34(0.10)
SeeSaw(Docea et al. 2024) 77.17(0.15) 81.44(0.10) 81.97(0.08) 74.42(0.15) 78.43(0.10) 80.01(0.08) 76.85(0.15) 79.83(0.10) 80.52(0.08)
Physics-Encoded GNN(Saleh et al. 2024) 76.48(0.14) 80.86(0.09) 81.72(0.12) 73.73(0.14) 77.85(0.09) 79.76(0.12) 76.16(0.14) 79.25(0.09) 80.27(0.12)
Transformer-KAN(Jiang et al. 2025) 78.82(0.12) 82.43(0.08) 83.36(0.07) 76.07(0.12) 79.42(0.08) 81.40(0.07) 78.50(0.12) 80.82(0.08) 81.91(0.07)
Diffeomorphic Deformation(Gong et al. 2024) 79.07(0.10) 83.30(0.11) 84.47(0.09) 76.32(0.10) 80.29(0.11) 82.51(0.09) 78.75(0.10) 81.69(0.11) 83.02(0.09)
PointNet(Qi et al. 2016) 81.25(0.22) 84.65(0.19) 85.05(0.15) 79.56(0.18) 82.66(0.16) 83.33(0.11) 83.25(0.17) 84.75(0.13) 85.21(0.09)
U-Net(Mendizabal, Márquez-Neila, and Cotin 2020) 80.56(0.10) 84.22(0.09) 84.88(0.11) 76.91(0.05) 79.83(0.08) 82.10(0.21) 78.10(0.12) 80.45(0.09) 81.33(0.14)
GraphGPS(Rampášek et al. 2022) 77.52(0.17) 80.63(0.13) 82.11(0.09) 73.72(0.15) 77.92(0.10) 80.66(0.16) 76.24(0.18) 79.31(0.15) 80.42(0.10)
SAB(Lee et al. 2019) 80.75(0.19) 84.44(0.07) 84.75(0.10) 78.88(0.11) 81.50(0.15) 82.83(0.16) 81.22(0.10) 83.01(0.08) 84.00(0.07)
@@
Table 3: Performance of various models on nonlinear simulations with individual graspers, combined graspers, and multitask learning. DCM (%) is reported with standard deviations. “Base” denotes the original model, “+Res.” augments it with a Kelvinlet-based residual, and “+Reg.” applies a Kelvinlet-inspired regularizer.